OpenStudy (anonymous):

Jacki is making a couple of stools for her patio. She needs 8 wooden legs that are each 3/5 yard long. She wants to cut them from redwood that is 5.25 yards long. How much wood will remain after Jacki cuts the 8 chair legs? Select one of the options below as your answer: A. 8.6 yards B. 4.8 yards C. 10.05 yards D. 0.45 yard

Directrix (directrix):

D. 0.45 yard ========== 5.25 yd - 8(3/5 yd) = 5.25 - 8(.6) = 5.25 - 4.8 = .45 yards

OpenStudy (anonymous):

cshalvey what are you talking about lol?

OpenStudy (anonymous):

Ahh, there was a answer there that I erased - you can disregard this, cheers ;)